# UDC: A Unified Neural Divide-and-Conquer Framework for Large-Scale Combinatorial Optimization Problems

### v1

@@ D.1 Hyperparameter
The hyperparameter of all 10 involved CO problems is listed in Table 7. The training of UDC on most CO problems adopts the varying-size [44] and varying-constraint [6] setting. ICAM and POMO used for conquering policy are 6-layers. DPN is 3-layers, and MatNet is 5-layer. AGNN for conquering sub-MIS is 12-layer with $d=256$. In training UDC for TSP, we don’t conduct the varying scale training in the first 50 epochs with $\alpha=80$.
@@
Two side conquering baseline TRUE TRUE FALSE TRUE
@@
Two side conquering baseline FALSE FALSE FALSE FALSE
@@ D.2 Experiment on OVRP, MIS, ATSP, KP, and min-max mTSP
The main part of this paper has evaluated the proposed UDC on large-scale instances of 5 representative CO problems. In this section. We will further evaluate the proposed UDC on more CO problems. OVRP is a variant of CVRP that has been widely involved in recent multi-task NCO work [71, 72]. We introduce this problem to test UDC’s ability to handle open-loop TSP based on CVRP. MIS is a classic CO problem widely used to evaluate heatmap-based solvers [35], which can test the capability of UDC to handle CO problems on graphs. We also study ATSP, and due to the poor performance of ordinary queuing policies on ATSP, we employ the MatNet [70] to solve sub-ATSP. The experiment on ATSP can evaluate the flexibility of the UDC framework and its ability to process matrix-based data. We also employ KP to test the ability of UDC to handle CO problems without neighborhood properties. Most importantly, min-max mTSP is introduced to test the ability of UDC to handle problems with different aggregation functions (i.e., the max aggregation function) except for min-sum (i.e., objective functions of all the else 9 CO problems).
@@
Methods Obj. Gap Time Obj. Gap Time Obj. Gap Time
LKH3 23.51 - 12 28.96 - 6s 39.88 - 6s
@@
Experiment on MIS: The MIS problem is a widely adopted CO problem for heatmap-based solvers. UDC can also solve the MIS problem by employing AGNN as the conquering policy. Following DIFUSCO, we evaluate the proposed UDC on the Erdős-Rényi (ER) [73] dataset. Large-scale instances are too dense so there may be situations where there are no variables to solve in the conquering stage, so UDC is not suitable for large-scale ER graphs (ER [11,000-12,000], p=0.02). We evaluate the MIS on ER[700-800] with p=0.15. We use the test set provided in [35], and the results on it are shown in Table9. UDC can also achieve good results on MIS after several conquering stages. Compared to the heatmap-based solver that originally relied on search algorithms, UDC shows better efficiency.
@@
Table 11: Objective function (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 500-node, 1,000-node (128 instances), and 2,000-node, 5,000-node KP (16 instances). OR-Tools can generate solutions in real time.
KP500,C=50 KP1,000,C=100
Method Obj. Gap Time Obj. Gap Time
@@
KP2,000,C=200 KP5,000,C=500
Method Obj. Gap Time Obj. Gap Time
@@
D.3 Experiments on Instances with Various Distributions.
The cross-distribution generalization ability of NCO solvers is necessary. So we evaluate the UDC on the Rotation distribution, and the Explosion distribution provided in Omni_VRP [75] on TSP. The result is shown in Table 13, UDC demonstrates outstanding robustness on large-scale CO problems with different distributions.
@@
TSP1,000, Rotation TSP1,000, Explosion CVRP1,000, Rotation CVRP1,000, Explosion
Method Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time Obj. (Gap) Time
Optimal 17.20 (0.00%) -- 15.63 (0.00%) -- 32.49 (0.00%) -- 32.31 (0.00%) --
POMO aug×$\times$8 24.58 (42.84%) 8.5m 22.70(45.24%) 8.5m 64.22 (97.64%) 10.2m 59.52 (84.24%) 11.0m
Omni_VRP+FS* 19.53(14.30%) 49.9m 17.75(13.38%) 49.9m 35.60 (10.26%) 56.8m 35.25 (10.45%) 56.8m
ELG aug×$\times$8 19.09(10.97%) 15.6m 17.37 (11.16%) 13.7m 37.04(14.00%) 16.3m 36.48(12.92%) 16.6m
UDC-$\boldsymbol{x}_{250}$($\alpha$=1) 18.19 (5.73%) 61s 16.76 (7.21%) 61s 34.91 (7.45%) 3.3m 34.75 (7.55%) 3.3m
D.4 Experiments on Very Large-scale
@@
However, on very large-scale TSP (results shown in Table 15), the UDC fails to converge to an outstanding performance in testing. We find that this is mainly due to a significant decrease in effectiveness when generalizing the dividing policy to 10,000-node TSP instances. So, we substitute the dividing policy to the random insertion used in [13], as shown in 16, The UDC with random-insertion-based initial solution achieved outstanding results. The experimental results preliminarily verify that in some problems such as TSP, UDC may generate unreliable dividing policies when processing large-scale generalization. In this case, using other initial solutions with a similar shape to the optimal solution can also obtain high-quality solutions.
@@
Methods Obj. Gap Time Obj. Gap Time
@@
Methods Obj. Gap Time Obj. Gap Time
@@
Methods Obj. Gap Time Obj. Gap Time
@@ Appendix E Ablation Study
Table 17: Objective function (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 500-node, 1,000-node, and 2,000-node instances.
@@
Methods Obj. Gap Time Obj. Gap Time Obj. Gap Time
@@
E.1 Ablation Study on $\alpha$ (Sampled Initia Solutions)
In this section, we demonstrate the different effectiveness of UDC related to the numbers of initial solutions (i.e., $\alpha$) for each designed CO problem. We use different $\alpha$ values to test the TSP, CVRP, OP, PCTSP, and SPCTSP (CO problems evaluated in the main paper) on their original test set. Table 17 exhibits the test results with different $\alpha$. Results with $\alpha=50$ always demonstrate significantly higher performance and much more time consumption.
However, the sensitivity of $\alpha$ on performance varies among different CO problems. In some problems, the differences between results with $\alpha=50$ and results with $\alpha=1$ appear to be more significant. In TSP, CVRP, and OP, the setting of $\alpha=50$ exhibits a significant effect, while in PCTSP and SPCTSP, using more initial solutions shows no significant effect. Moreover, On 2,000-node problems TSP, CVRP, and OP, the effect of $\alpha=50$ is particularly pronounced. Since the difference in initial solutions is entirely due to the sampling of dividing policy, a more significant distinction between $\alpha=50$ and $\alpha=1$ variants might mean that the initial solution prediction is more difficult in these instances. Meanwhile, the more significant difference in 2,000-node instances indicates that compared to (S)PCTSP, the dividing policy has worse scaling ability on TSP, CVRP, and OP.
E.2 Ablation Study on $r$ (Convergence Speed In Testing)
The proposed UDC adopts a multi-conquering stage ($r$ stages) as common variants. Therefore, the effectiveness of our testing phase is closely related to the testing time. We record the iterative convergence curves of the objective function value (i.e., Obj.) during testing TSP1,000, CVRP1,000, and PCTSP1,000 with UDC($\alpha$=1), UDC($\alpha$=50), GLOP [13], LEHD (RRC) [8], and DIFUSCO [39]. The results are shown in Figure 6, and curves are generated under the same test set and similar computation resources. We can obtain the following conclusions:
1. For the two variants of UDC, UDC($\alpha$=1) can generate better solutions compared to UDC($\alpha$=50) in a limited period but more quickly traps in local optima and stops optimizing. For PCTSP, there is no significant difference in the final convergence result between the two versions.
2. Compared to the advanced neural divide-and-conquer method GLOP, the objective functions of both the two UDC variants significantly lead when going through the same testing time, and the final convergence results of UDC variants are better.
(a) TSP1,000
(b) CVRP1,000
(c) PCTSP1,000
Figure 5: Testing curves for ablation study.
E.3 Ablation Study on Components
@@
Moreover, we also evaluate the significance of constraint normalization on SPCTSP and PCTSP. As shown in Figure 5(c), The final performance after 250 epochs and training efficiency of UDC is strongly nerfed without the normalization on the total prize lower-bound in sub-(S)PCTSPs.

### v2

@@ D.1 Hyperparameter
Hyperparameters of all 10 involved CO problems are listed in Table 7. Training UDC on most CO problems adopts the varying-size [44] and varying-constraint [6] setting. ICAM and POMO used for conquering policy are 6-layers. DPN is 3-layers, and MatNet is 5-layer. AGNN for conquering sub-MIS is 12-layer with $d=256$. In training UDC for TSP, we don’t conduct the varying scale training in the first 50 epochs with $\alpha=80$.
@@
Two sides conquering baseline TRUE TRUE FALSE TRUE
@@
Two sides conquering baseline FALSE FALSE FALSE FALSE
@@
It can be seen that UDC adopts similar setups for solving different problems. The third to last and the second to last rows of Table 7 are related to the environment and the last row unable to use a consistent setup and some flexibility in choices. Ablation studies in Section 5.2 and Appendix E discuss the setting of some listed hyperparameters, such as $\alpha$ in testing, DCR enable=False for KP, and the selection of conquering Policy in KP and TSP.
@@ D.2 Experiment on OVRP, MIS, ATSP, KP, and min-max mTSP
The main part of this paper has evaluated the proposed UDC on large-scale instances of 5 representative CO problems. In this section. We will further evaluate the proposed UDC on more CO problems. OVRP is a variant of CVRP that has been widely involved in recent multi-task NCO work [75, 76]. We introduce this problem to test UDC’s ability to handle open loops and distant point-to-point relationships based on CVRP. MIS is a classic CO problem widely used to evaluate heatmap-based solvers [35], which can test the capability of UDC to handle CO problems on graphs. The experiment on ATSP can evaluate the flexibility of the UDC framework and its ability to process matrix-based data. We also employ KP to test the ability of UDC to handle CO problems without neighborhood properties. Most importantly, min-max mTSP is introduced to test the ability of UDC to handle problems with different aggregation functions (i.e., the max aggregation function) except for min-sum (i.e., objective functions of all the else 9 CO problems). For OVRP, KP, ATSP, and min-max mTSP in this section, UDC works to minimize rather than maximize (like OP and MIS) their objective functions.
@@
Method Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time
LKH3 23.51 - 16m 28.96 - 32m 39.88 - 27m
@@
Experiment on MIS: The MIS problem is a widely adopted CO problem for heatmap-based solvers. UDC can also solve the MIS problem by employing AGNN as the conquering policy. Following DIFUSCO, we evaluate the proposed UDC on the Erdős-Rényi (ER) [77] dataset. We evaluate the MIS on ER[700-800] with $p$=0.15. We use the test set provided in [35], and the results on it are shown in Table9. UDC can also achieve good results on MIS after several conquering stages. Compared to the heatmap-based solver that originally relied on search algorithms, UDC shows better efficiency.
@@
Table 11: Objective function (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 500-node, 1,000-node (128 instances), and 2,000-node, 5,000-node KP (16 instances). OR-Tools can generate solutions in real-time.
KP500,$W$=50 KP1,000,$W$=100
Method Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time
@@
KP2,000,$W$=200 KP5,000,$W$=500
Method Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time
@@
D.3 Experiments on TSP & CVRP Instances with Various Distributions
The cross-distribution generalization ability of NCO solvers is also necessary. So we evaluate the UDC on the Rotation distribution, and the Explosion distribution provided in Omni_VRP [79] on TSP and CVRP. The result is shown in Table 13, UDC demonstrates outstanding robustness on large-scale CO problems with different distributions.
@@ D.3 Experiments on TSP & CVRP Instances with Various Distributions
TSP1,000, Rotation TSP1,000, Explosion
Method Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time
Optimal 17.2 0.00% - 15.63 0.00% -
POMO 24.58 42.84% 8.5m 22.7 45.24% 8.5m
Omni_VRP+FS* 19.53 14.30% 49.9m 17.75 13.38% 49.9m
ELG 19.09 10.97% 15.6m 17.37 11.16% 13.7m
UDC-$\boldsymbol{x}_{250}$($\alpha$=1) 18.19 5.73% 61s 16.76 7.21% 61s
CVRP1,000, Rotation CVRP1,000, Explosion
Method Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time
Optimal 32.49 0.00% - 32.31 0.00% -
POMO 64.22 97.64% 10.2m 59.52 84.24% 11.0m
Omni_VRP+FS* 35.6 10.26% 56.8m 35.25 10.45% 56.8m
ELG 37.04 14.00% 16.3m 36.48 12.92% 16.6m
UDC-$\boldsymbol{x}_{250}$($\alpha$=1) 34.91 7.45% 3.3m 34.75 7.55% 3.3m
D.4 Experiments on MIS Instances with Various Distributions
In addition to the TSP and CVRP, we conduct generalization experiments on the MIS problems. We conduct zero-shot generalization on the model trained by ER graphs with $N\sim U(700,800)$ and $p$=0.15 to five other datasets, including:
  • A random ER graph dataset with $N$=720 and $p$=0.05. (average 12,960 undirected edges)
  • A random ER graph dataset with $N$=720 and $p$=0.25. (average 64,800 undirected edges)
  • A random ER graph dataset with $N$=2,000 and $p$=0.05. (average 100,000 undirected edges)
  • The ego-Facebook data in Stanford Large Network Dataset Collection (SNAP) (https://snap.stanford.edu/data/). (4,039 nodes, 88,234 undirected edges)
  • The feather-lastfm-social data in SNAP. (7,624 nodes, 27,806 undirected edges)
All the first three ER random datasets contain 128 graph instances, and we consider the last two data as a single graph. The table below exhibits the comparison results between UDC and the SOTA heuristics KaMIS (implemented on https://github.com/KarlsruheMIS/KaMIS) on the original in-domain dataset (ER-[700-800], $p$=0.15) and the five out-of-domain datasets mentioned above. Obj. represents the objective function value, Gap represents the performance gap compared to the best algorithm and Time is the time consumption for solving the whole dataset.
Dataset ER, $N$=720, $p$=0.05 ER-[700-800], $p$=0.15(in-domain)
Method Obj.$\uparrow$ Gap Time Obj.$\uparrow$ Gap Time
KaMIS 99.5 - 40.6m 44.9 - 52.13m
UDC-$x_{0}$($\alpha=50$) 79.5 20.10% 8s 41.0 8.62% 40s
UDC-$x_{50}$($\alpha=50$) 89.5 10.07% 9.03m 42.9 4.44% 21.05m
UDC-$x_{250}$($\alpha=50$) 94.5 5.03% 44.84m 43.8 2.41% 1.73h
Dataset ER, $N$=720, $p$=0.25 ER, $N$=2,000, $p$=0.05
Method Obj.$\uparrow$ Gap Time Obj.$\uparrow$ Gap Time
KaMIS 28.2 - 1.4h 133.9 - 3h
UDC-$x_{0}$($\alpha=20$) 21.6 23.46% 58s 116.9 12.72% 2m
UDC-$x_{50}$($\alpha=20$) 25.3 10.54% 21m 119.1 11.05% 21m
UDC-$x_{250}$($\alpha=20$) 26.4 6.59% 63m 122.9 8.25% 90m
Dataset SNAP-ego-Facebook SNAP-feather-lastfm-social
Method Obj.$\uparrow$ Gap Time Obj.$\uparrow$ Gap Time
KaMIS 1052.0 - 155s 4177.0 - 0.1s
UDC-$x_{0}$($\alpha=1$) 767.0 27.09% 6s 3622.0 13.29% 2s
UDC-$x_{250}$($\alpha=1$) 901.0 14.35% 11s 3751.0 10.20% 13s
UDC-$x_{2500}$($\alpha=1$) 1009.0 4.09% 60s 4067.0 2.63% 80s
Compared to KaMIS, UDC variants exhibit relatively stable performance gaps across datasets with various node sizes (n), sparsity (p), and generation methods, demonstrating good generalization ability. Moreover, except for the extremely sparse graph SNAP-feather-lastfm-social, UDC variants have significant advantages over KaMIS in terms of time.
Unfortunately, we are unable to re-implement any learning-based methods for MIS (e.g. Difusco [39] and T2T [65]), so we are unable to compare the generalization ability with other learning based algorithms. However, the stable performance gaps on different datasets partially demonstrate certain out-of-domain generalization performance of UDC on MIS.
D.5 Experiments on Very Large-scale
@@
However, on very large-scale TSP (results shown in Table 16), the UDC fails to converge to an outstanding performance in testing. We find that this is mainly due to a significant decrease in effectiveness when generalizing the dividing policy to 10,000-node TSP instances. So, we substitute the dividing policy to the random insertion used in [13], as shown in Table 17, The UDC with random-insertion-based initial solution achieved outstanding results. The experimental results preliminarily verify that in some problems such as TSP, UDC may generate unreliable dividing policies when processing large-scale generalization. In this case, using other initial solutions with a similar shape to the optimal solution can also obtain high-quality solutions.
@@
Methods Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time
@@
Methods Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time
@@
Methods Obj.$\downarrow$ Gap Time Obj.$\downarrow$ Gap Time
@@ Appendix E Ablation Study
E.1 Ablation Study on $r$ (Convergence Speed In Testing)
The proposed UDC adopts the multi-conquering stage ($r$ stages) in common variants. Therefore, as an iterative method, the effectiveness of our testing phase is closely related to the testing time. We record the iterative convergence curves of the objective function value (i.e., Obj.) during testing TSP1,000, CVRP1,000, and PCTSP1,000 with UDC($\alpha$=1), UDC($\alpha$=50), GLOP [13], LEHD (RRC) [8], and DIFUSCO [39]. The results are shown in Figure 5, and curves (except for DIFUSCO, which uses reported results) are generated under the same test set and computation resources. We can obtain the following conclusions:
1. For the two variants of UDC, UDC($\alpha$=1) can generate better solutions compared to UDC($\alpha$=50) in a limited period but will trap into local optima quickly. For PCTSP, there is no significant difference in the final convergence result between the two versions.
2. Compared to the advanced neural divide-and-conquer method GLOP, the objective functions of both the two UDC variants always significantly lead when going through the same testing time.
(a) TSP1,000
(b) CVRP1,000
(c) PCTSP1,000
Figure 5: Testing curves for ablation study.
E.2 Ablation Study on $\alpha$ (Sampled Initial Solutions)
In this subsection, we demonstrate the change in effectiveness of UDC relating to the numbers of initial solutions (i.e., $\alpha$) for each CO problem. We use different $\alpha$ values to test the TSP, CVRP, OP, PCTSP, and SPCTSP (CO problems evaluated in the main paper) on their original test set. As exhibited in Table 18, results with $\alpha=50$ always demonstrate significantly better performance but much more time consumption.
However, the sensitivity of $\alpha$ on performance varies among different CO problems. In TSP, CVRP, and OP, the differences between results with $\alpha=50$ and results with $\alpha=1$ appear to be quite significant, while in PCTSP and SPCTSP, employing more initial solutions demonstrates no significant effect. Moreover, On 2,000-node problems TSP, CVRP, and OP, the effect of $\alpha=50$ is particularly pronounced. Since the difference in initial solutions is entirely from sampling the dividing policy, a more significant distinction between $\alpha=50$ and $\alpha=1$ variants might mean that the initial solution prediction is more difficult in these instances. Meanwhile, the more significant difference in 2,000-node instances indicates that compared to (S)PCTSP, the dividing policy has worse scaling ability on TSP, CVRP, and OP.
Table 18: Ablation on the number of initial solution $r$. Objective function values (i.e., Obj.), Gap to the best algorithm (i.e., Gap) on 500-node, 1,000-node, and 2,000-node instances are listed.
@@ E.2 Ablation Study on $\alpha$ (Sampled Initial Solutions)
Methods Obj.$\uparrow$ Gap Time Obj.$\uparrow$ Gap Time Obj.$\uparrow$ Gap Time
@@
E.3 Ablation Study on the Generation Method of Initial Solutions
To check the ability of the dividing policy, we further conduct experiments on TSP with different generation methods of initial solutions, and the results shown in Table 19 demonstrate the significance of a good initial solution to the final performance. The UDC variants with a random initial solution or a nearest greedy initial solution cannot converge to good objective values even with 50 conquering stages. Moreover, from TSP500 to TSP2,000, the dividing policy of UDC produces similar results to the heuristic algorithm random insertion (used in GLOP [13]), which verifies the quality of the dividing policy.
Table 19: Ablation on different algorithms as initial solutions $\boldsymbol{x}_{0}$. The conquering stages of all the variants are the constructive of UDC. Random-UDC directly represents a random solution as the initial solution. Nearest Greedy-UDC uses the nearest greedy algorithm (starting from the first node) and Random Insertion-UDC employs the random insertion heuristic as the initial solution. $\alpha=1$ in all UDC variants. UDC-$\boldsymbol{x}_{2}$ and UDC-$\boldsymbol{x}_{50}$ is the original version. Obj. is the objective value and Gap represents the gap to the best method.
Method TSP500 TSP1,000 TSP2,000
Method Obj. Gap Obj. Gap Obj. Gap
Concorde 16.52 - 23.12 - 32.45 -
Random-UDC-$\boldsymbol{x}_{2}$ 33.98 105.68% 67.82 193.34% 134.98 315.98%
Random-UDC-$\boldsymbol{x}_{50}$ 26.11 58.01% 52.48 126.98% 104.48 221.96%
Nearest Greedy-UDC-$\boldsymbol{x}_{2}$ 18.48 11.87% 26.42 14.27% 37.48 15.50%
Nearest Greedy-UDC-$\boldsymbol{x}_{50}$ 17.88 8.23% 25.75 11.38% 36.80 13.40%
Random Insertion-UDC-$\boldsymbol{x}_{2}$ 17.09 3.45% 24.08 4.17% 34.00 4.78%
Random Insertion-UDC-$\boldsymbol{x}_{50}$ 16.84 1.93% 23.75 2.73% 33.55 3.38%
UDC-$\boldsymbol{x}_{2}$($\alpha=1$) 17.12 3.62% 24.06 4.05% 34.44 6.13%
UDC-$\boldsymbol{x}_{50}$($\alpha=1$) 16.92 2.41% 23.68 2.42% 33.60 3.54%
E.4 Ablation Study on Components
@@
Moreover, we also evaluate the significance of constraint normalization on PCTSP. As shown in Figure 6(c), The final performance after 250 epochs and training efficiency of UDC is strongly nerfed without the normalization on the total prize lower-bound in sub-(S)PCTSPs.
@@
E.5 Ablation Study on Disable DCR in KP
For KP, there will be no sub-optimal connection as shown in Figure 2, so there is no logical difference in whether DCR is enabled or not. We provide the KP results of Enable DCR in Table 20 and the results support the above conclusion.
Table 20: Ablation on whether to enable the DCR in training UDC for KP. The version with disabling DCR (i.e., Disable DCR) is the same as Table 11 in the original paper and the datasets are also the same. The value shown in the table is the gap to optimal (OR-Tools).
Method KP500 KP1,000 KP2,000 KP5,000
UDC-$\boldsymbol{x}_{50}$(Disable DCR) 0.0247% 0.0260% 0.0237% 0.0249%
UDC-$\boldsymbol{x}_{250}$(Disable DCR) 0.0083% 0.0106% 0.0105% 0.0100%
UDC-$\boldsymbol{x}_{50}$(Enable DCR) 0.0105% 0.0258% 0.0254% 0.0273%
UDC-$\boldsymbol{x}_{250}$(Enable DCR) 0.0080% 0.0117% 0.0128% 0.0095%
E.6 Ablation Study on Conquering Policies
UDC exhibits flexibility in choosing a constructive solver for conquering policy. For well-developed CO problems like TSP and KP, there are multiple constructive solvers available. Section 5.2 presents an ablation study on the selection of these constructive solvers for TSP. In this subsection, we conducted a similar ablation study on KP, and the results are displayed in Table 21. Both sets of results indicate that the final performance of UDC is not sensitive to the model selection for the conquering policy.
Therefore, when generalizing UDC to new CO problems, it may not be necessary to specifically choose which available constructive solver to model the conquering policy.
Table 21: Ablation on the choice of constructive model as the conquering policy in KP. ICAM and POMO are two available constructive solvers for KP. The datasets are the same as Table 11 in the original paper. The value shown in the table is the gap to optimal (OR-Tools).
Method KP500 KP1,000 KP2,000 KP5,000
UDC-$\boldsymbol{x}_{50}$(ICAM) 0.0247% 0.0260% 0.0237% 0.0249%
UDC-$\boldsymbol{x}_{250}$(ICAM) 0.0083% 0.0106% 0.0105% 0.0100%
UDC-$\boldsymbol{x}_{50}$(POMO) 0.0255% 0.0252% 0.0253% 0.0242%
UDC-$\boldsymbol{x}_{250}$(POMO) 0.0098% 0.0107% 0.0119% 0.0113%